# Conceptualization; Methodology; Resources; Investigation; Writing original draft; Writing, review, and editing

### v1

@@ Abstract
Automatic speech recognition (ASR) plays a pivotal role in our daily lives, offering utility not only for interacting with machines but also for facilitating communication for individuals with either partial or profound hearing impairments. The process involves receiving the speech signal in analogue form, followed by various signal processing algorithms to make it compatible with devices of limited capacity, such as cochlear implants (CIs). Unfortunately, these implants, equipped with a finite number of electrodes, often result in speech distortion during synthesis. Despite efforts by researchers to enhance received speech quality using various state-of-the-art signal processing techniques, challenges persist, especially in scenarios involving multiple sources of speech, environmental noise, and other circumstances. The advent of new artificial intelligence (AI) methods has ushered in cutting-edge strategies to address the limitations and difficulties associated with traditional signal processing techniques dedicated to CIs. This review aims to comprehensively review advancements in CI-based ASR and speech enhancement, among other related aspects. The primary objective is to provide a thorough overview of metrics and datasets, exploring the capabilities of AI algorithms in this biomedical field, summarizing and commenting on the best results obtained. Additionally, the review will delve into potential applications and suggest future directions to bridge existing research gaps in this domain.
keywords:
@@ 1 Introduction
ASR is pivotal in authentication systems, safeguarding the security and privacy of sensitive information. The integrity of audio speech can be verified through techniques-based ASR such as adversarial attacks detections [3], steganalysis [4, 5, 6], speech biometric [7] and more. Beyond the realm of communication, ASR finds itself at the heart of various applications, each playing a unique role in different domains. Speaker recognition, a facet of ASR, is not merely confined to enhancing security measures. It has evolved into a versatile tool employed in healthcare, where the identification of individuals through their unique vocal signatures holds promise for personalized patient care. This is particularly relevant in scenarios where quick and secure authentication is crucial, such as accessing medical records or authorizing medical procedures. Event recognition, another dimension of ASR, is a game-changer in sectors ranging from security to healthcare. In the former, ASR algorithms analyze audio data to automatically detect and categorize specific events, reinforcing surveillance capabilities. In healthcare, event recognition becomes a powerful tool for monitoring and early detection of health-related events. For instance, in the context of cardiac health, ASR can aid in identifying anomalies in heart sounds, potentially enabling early intervention and preventive measures [8]. Source separation, the ability to discern and isolate individual sound sources from complex audio signals, is a boon in fields like entertainment and music production. However, its significance extends into the realm of biomedical research, where ASR plays a pivotal role in decoding the intricate language of physiological signals. In the context of cochlear implants, source separation becomes a critical component in enhancing the auditory experience for individuals with hearing impairments.
[…]
1.1 Related work
@@
Many reviews have been written in the context of CI. For example, [9] discussed the advantages offered by machine learning (ML) to cochlear implantation, such as analyzing data to personalize treatment strategies. It enhances accuracy in speech processing optimization, surgical anatomy location prediction, and electrode placement discrimination. Besides, it delves into its applications, including optimizing cochlear implant fitting, predicting patient threshold levels, and automating image-guided CI surgery. The review discusses some open novel opportunities for research, emphasizing the need for high-quality data inputs and addressing concerns about algorithm transparency in clinical decision-making for improved patient care. Similarly, the review written by Manero et al. [10] details some benefits of employing artificial intelligent (AI) in enhancing CI technology, involving adaptive sound processing, acoustic scene classification, and auditory scene analysis. The authors discuss AI-driven advancements aiming to optimize sound signals, adapt to diverse environments, and improve speech perception for individuals with hearing loss, ultimately enhancing their overall quality of life. Also, the review [11] explores three main topics: direct-speech neuroprosthesis, which involves decoding speech from the sensorimotor cortex using AI, including the synthesis of produced speech from brain activity. The second topic is a top-down exploration of pediatric cochlear implantation using ML, delving into the applications of ML in pediatric cochlear implantation. Lastly, the paper considers the potential of AI to solve the hearing-in-noise problem, examining its capabilities in addressing challenges related to hearing in noisy environments. Moreover, the review [12] critically examines the current landscape of tele-audiology practices, highlighting both their constraints and potential opportunities. Specifically, it explores intervention and rehabilitation efforts for CIs, focusing on remote programming and the concept of self-fitting CIs. Recently, a review written by Henry et al. in 2023 [13] conducts a comprehensive review of noise reduction algorithms employed in CIs. Maintaining a general classification based on the number of microphones used—single or multiple channels—the analysis extends to incorporate recent studies showcasing a growing interest in ML techniques. The review culminates with an exploration of potential research avenues that hold promise for future advancements in the field. Table 1 offers a comparative analysis of the proposed review in contrast to other discussed AI-based CI reviews and surveys.
@@
Table 1: Comparison of this review against other existing AI-based CI reviews and surveys. Tick marks (✓) indicate that a particular field has been considered, while cross marks () indicate that a field has been left unaddressed. The symbol (★) indicates that the most critical concerns of a field have not been addressed.
Ref.
Year
Description
Backg. TCI Metrics datasets MLT DLT Apps RGCs FDs
[9]
2019
The intersection of ML and CIs
[10]
2022
Explores materials and devices designed for medical care
[11]
2022
AI in otolaryngology and communication sciences
[12]
2022
Remote audiology services
[13]
2023
Signal processing in CI for noise reduction
Our
2024
Advanced AI algorithms for CIs
@@
1.2 Statistics on investigated papers
Recently, there has been a surge in publications related to AI-based CI. The review methodology entails defining the search strategy and study selection criteria. Criteria for inclusion, like keyword relevance and impact, shape the quality assessment protocol. A comprehensive search was conducted on databases such as Scopus and Web of Science. Keywords were extracted for theme clustering, resulting in a formulated query to gather advanced AI-based CI studies. The research query fetches references from papers containing the keywords "Cochlear implant" or "Hearing loss" and "Artificial intelligence" in their abstracts, titles, or authors’ keywords. It subsequently refines these papers, focusing on those that also include "Machine learning," "Deep learning," or "Reinforcement learning." Figure 1 illustrates the most frequently used keywords by the authors in the titles, abstracts, and keywords of the selected papers.
@@ 1.3 Motivation and contribution
  • The implementation of CI, along with a comprehensive elucidation of the taxonomy encompassing ML and DL-based CI, is thoroughly expounded upon. Additionally, recommended frameworks for AI-based CI are thoroughly discussed and succinctly summarized in tables for enhanced clarity.
@@
  • Delve into the existing gaps in AI-driven CI, offer insights, and propose novel ideas to address these gaps. Additionally, explore potential avenues for future research to deepen comprehension and provide valuable guidance for subsequent investigations.
The subsequent sections are organized as follows: Section 2 delves into the background in speech processing for, outlining datasets and metrics. Section 3 discusses the methodology employed for CI based on AI. Section 4 presents the medical applications and impact of applying AI on CI. Section 5 offers a comprehensive discussion research gaps, future directions, and perspectives. Finally, Section 6 concludes the paper with implications and future research directions.
2 Background in speech processing for CIs
2.1 Speech processing for CIs
CIs are electronic devices that can be implanted in one or both ears to restore some level of hearing for individuals with partial or severe deafness. CIs comprise an external part with a microphone and speech processor and an internal part with a receiver-stimulator and electrode array as shown in Figure 3. They convert sounds into electrical signals, stimulating the auditory nerve to enable sound perception in individuals with profound hearing loss [14].
@@ 2.1 Speech processing for CIs
The incoming sound is divided into multiple frequency channels using bandpass filters and then processed by envelope detectors. Non-linear compressors adjust the dynamic range of the envelope for each patient. The compressed envelope amplitudes are then utilized to modulate a fixed-rate biphasic carrier signal. A current source converts voltage into pulse trains of current, which are delivered to electrodes placed along the cochlea in a non-overlapping manner. This stimulation method is called continuous interleaved sampling (CIS). Another coding strategy, known as advanced combination encoder (ACE), uses a greater number of channels and dynamically selects the "n-of-m" bands with the largest envelope amplitudes (prior to compression). Only the corresponding "n" electrodes are stimulated. A popular device widely used for CIs, such as the cochlear nucleus, typically has 22 channels. The sound processor, which usually contains a microphone, battery, and other components, can be worn either behind-the-ear (BTE) or off-the-ear (OTE). A headpiece holds a transmitter coil, positioned externally above the ear, while internally, a receiver coil, stimulator, and electrode array are implanted. The SP includes a digital signal processor (DSP) with memory units (maps) that store patient-specific information. An audiologist configures these maps during the fitting process, adjusting thresholds for each electrode, including T-Levels (the softest current levels audible to the CI user) and C/M-Levels (current levels perceived as comfortably loud), as well as the stimulation rate or programming strategy. Data (pulse amplitude, pulse duration, pulse gap, etc.) and power are sent through the skull via an radio frequency signal from the transmitter coil to the receiver coil. The stimulator decodes the received bitstream and converts it into electric currents to be delivered to the cochlear electrodes. High-frequency signals stimulate electrodes near the base of the cochlea, while low-frequency signals stimulate electrodes near the apex. The CI stimulates the auditory nerve afferents, which connect to the central auditory pathways. However, compared to individuals with normal hearing (NH), CI users face more difficulties in speech perception, particularly in noisy environments. Hearing loss can be caused by various factors, including natural aging, genetic predisposition, exposure to loud sounds, and medical treatments. Damage to the hair cells in the inner ear often leads to a reduced dynamic range of hearing, as well as decreased frequency selectivity and discriminative ability in speech processing. To evaluate the effectiveness of CIs in speech perception amid noise, listening tests involving both normal hearing individuals and CI users are commonly conducted. These tests typically employ a combination of speech utterances from a recognized speech corpus and background noises like speech-weighted noise and babble. Alternatively, vocoder simulations can be utilized alongside speech intelligibility metrics. While sentence-based tests are frequently employed, other stimuli such as vowels, consonants, and phonemes are also used. As a result, noise reduction techniques are increasingly employed to enhance the performance of CIs in challenging environments [13].
2.2 Datasets
@@
Researchers have utilized numerous datasets to validate their proposed schemes, comprising both widely recognized publicly available sets and locally generated ones. These datasets fall into two categories: speech or images. Table 2 provides a summary of these datasets, detailing their characteristics, citing studies that have utilized them, and indicating their availability through links or references.
Table 2: List of publicly available datasets used for CIs applications
@@
PhonDat 1 21587 201 It is from the Bavarian Archive For Speech Signals, contains 21.4 hours of speech data, and it is orthographically transcribed and phonemically annotated. [19],[20] Link [^1] https://www.phonetik.uni-muenchen.de/Bas/BasPD1eng.html TIMIT 6300 630 The dataset consists of phonemically and lexically transcribed speech from American English speakers belonging to diverse demographics and dialects. It provides comprehensive information with time-aligned orthographic, phonetic, and word transcriptions. Additionally, each utterance is accompanied by its corresponding 16-bit, 16kHz speech waveform file, ensuring a complete and detailed dataset for analysis and experimentation in ASR and acoustic-phonetic studies. [21],[22],[23] Link [^2] https://www.kaggle.com/mfekadu/darpa-timit-acousticphonetic-continuous-speech GSC 18 Hours 30 The dataset was gathered using crowd-sourcing. It consists of 65,000 recordings, each lasting one second, and contains 30 brief words. Among these words, 20 commonly used ones were spoken five times by the majority of participants, while 10 other words (considered unfamiliar) were spoken only once. [24] Link [^3] https://research.googleblog.com/2017/08/launching-speech-commands-dataset.html WSJ0-2mix - - This dataset is a collection of mixed speech recordings that are used for speech recognition. It utilizes utterances from the Wall Street Journal (WSJ0) corpus to create these mixtures. [25] Link  [^4] https://www.merl.com/research/highlights/deep-clustering LibriVox 547 Hours - The LibriVox dataset consists of sentence-aligned German audio, text, and English translations from audio books. [22] Link  [^5] https://librivox.org/ HSM 20 30 The development of the German HSM Sentence Test aimed to provide an ample amount of test sentences for the repeated assessment of speech comprehension among CI users. [22],[26] [27] CEC1 6000 24 The dataset features simulated living rooms with static sources, including a single target speaker, interferer (competing talker or noise), and a large target speech database of English sentences produced by 40 British English speakers. [26] Link [^6] https://claritychallenge.org/clarity_CEC1_doc/docs/cec1_data DEMAND 560 6 The dataset consists of 15 recordings capturing acoustic noise in various environments. These recordings were made using a 16-channel array, with microphone distances ranging from 5 cm to 21.8 cm. [22] Link [^7] https://www.kaggle.com/datasets/chrisfilo/demand THCHS-30 35 Hours 50 The dataset is a free Chinese speech corpus accompanied by resources such as lexicon and language models. [28] Link [^8] https://www.openslr.org/18/ BCP 55938 20 The bern cocktail party (BCP) dataset contains Cocktail Party scenarios with individuals wearing CI audio processors and a head and torso simulator. Recorded in an acoustic chamber, it includes multi-channel audio, image recordings, and digitized microphone positions for each participant. [29] [30] ikala 252-30s 206 The dataset comprises audio recordings consisting of vocal and backing track music with a sampling rate of 44100 Hz. Each music track is a stereo recording, where one channel contains the singing voice and the other channel contains the background music. All tracks were performed by professional musicians and featured a group of six singers, evenly split between three females and three males. [31] Link [^9] https://zenodo.org/records/3532214 MUSDB 150 4 The dataset is a collection of music tracks specifically designed for music source separation research. It consists of professionally mixed songs across various genres, with individual tracks isolated for vocals, drums, bass, and other accompaniment. [31] Link [^10] https://sigsep.github.io/datasets/musdb.html CQ500 491 - The dataset includes anonymized dicoms files, along with the interpretations provided by radiologists. The interpretations were conducted by three radiologists who have 8, 12, and 20 years of experience in interpreting cranial CT scans, respectively. [32] Link [^11] http://headctstudy.qure.ai/dataset ECochG 21 - The dataset captures inner ear potentials in response to acoustic stimulation, specifically focusing on cochlear implant recipients. [33],[34],[35] Link [^12] https://zenodo.org/records/7092661
@@
2.3 Metrics
Multiple evaluation metrics are utilized during the training and validation of any DL models, including AI-based CI. These metrics, integral to the confusion matrix, are widely known and applicable across various data types such as speech or image. They include Accuracy (Acc), Sensitivity (Sen), Recall (Rec), Specificity (Spe), Precision (Pre), F1 score (F1), and receiver operating characteristic curve (ROC). Moreover, different metrics play roles in prediction task. For instance, intersection over union (IoU) assesses overlap, and mean absolute error (MAE) quantifies absolute differences.For a comprehensive understanding of the metrics discussed, including their equations, refer to the details provided in [5, 36]. Other metrics that are widely used for CI are summarized in Table LABEL:tab:metrics.
Metric
Formula
Description
DCS
$\frac{{2\cdot|A\cap B|}}{{|A|+|B|}}$
The metric dice coefficient similarity (DCS) is used to evaluate the performance of the vestibule segmentation network. The Dice coefficient is a widely used similarity metric in image segmentation tasks. It measures the overlap between the predicted segmentation mask and the ground truth mask.
ASD
$\frac{\sum_{a\in S(A)}{\min_{b\in S(B)}(a-b)}+\sum_{b\in S(B)}{\min_{a\in S(A)}(b-a)}}{|S(A)|+|S(B)|}$

### v3

@@ Abstract
Automatic speech recognition (ASR) plays a pivotal role in our daily lives, offering utility not only for interacting with machines but also for facilitating communication for individuals with partial or profound hearing impairments. The process involves receiving the speech signal in analog form, followed by various signal processing algorithms to make it compatible with devices of limited capacities, such as cochlear implants (CIs). Unfortunately, these implants, equipped with a finite number of electrodes, often result in speech distortion during synthesis. Despite efforts by researchers to enhance received speech quality using various state-of-the-art (SOTA) signal processing techniques, challenges persist, especially in scenarios involving multiple sources of speech, environmental noise, and other adverse conditions. The advent of new artificial intelligence (AI) methods has ushered in cutting-edge strategies to address the limitations and difficulties associated with traditional signal processing techniques dedicated to CIs. This review aims to comprehensively cover advancements in CI-based ASR and speech enhancement, among other related aspects. The primary objective is to provide a thorough overview of metrics and datasets, exploring the capabilities of AI algorithms in this biomedical field, and summarizing and commenting on the best results obtained. Additionally, the review will delve into potential applications and suggest future directions to bridge existing research gaps in this domain.
Index Terms:
Automatic speech recognition, Cochlear implant, Deep learning, Profound hearing loss, Speech enhancement, Machine learning
@@
ESTOI
extended short-time objective intelligibility
Bi-LSTM
bidirectional long short term memory
AAD
auditory attention decoding
SVM
support vector machine
@@ I Introduction
ASR is pivotal in authentication systems, safeguarding the security and privacy of sensitive information. The integrity of audio speech can be verified through techniques-based ASR such as adversarial attack detection [3], steganalysis [4, 5, 6], speech biometrics [7], and more. Beyond the realm of communication, ASR finds itself at the heart of various applications, each playing a unique role in different domains. Speaker recognition, a facet of ASR, is not merely confined to enhancing security measures. It has evolved into a versatile tool employed in healthcare, where the identification of individuals through their unique vocal signatures holds promise for personalized patient care. This is particularly relevant in scenarios where quick and secure authentication is crucial, such as accessing medical records or authorizing medical procedures. Event recognition, another dimension of ASR, is a game-changer in sectors ranging from security to healthcare. In the former, ASR algorithms analyze audio data to automatically detect and categorize specific events, reinforcing surveillance capabilities. In healthcare, event recognition becomes a powerful tool for monitoring and early detection of health-related events, recognizing speech from noisy environment [8], understanding person with dysarthria severity [9], and more. In the context of cardiac health, ASR can aid in identifying anomalies in heart sounds, potentially enabling early intervention and preventive measures [10]. Source separation, the ability to discern and isolate individual sound sources from complex audio signals, is a boon in fields like entertainment and music production. However, its significance extends into the realm of biomedical research, where ASR plays a pivotal role in decoding the intricate language of physiological signals. In the context of cochlear implants, source separation becomes a critical component in enhancing the auditory experience for individuals with hearing impairments.
The CIs, designed to restore hearing in individuals with severe hearing loss or deafness, rely on ASR for optimizing their functionality. ASR contributes significantly to the improvement of speech perception in CI users by enhancing the processing and interpretation of auditory signals. CIs work by converting sound waves into electrical signals that stimulate the auditory nerve, bypassing damaged parts of the inner ear. ASR complements this process by aiding in the recognition and translation of spoken language. The technology plays a crucial role in optimizing speech understanding for CI users by refining the interpretation of varied speech patterns, tones, and nuances. Moreover, ASR in the context of CIs extends beyond basic speech recognition. It contributes to the recognition of environmental sounds, facilitating a more immersive auditory experience for individuals with hearing impairments. This is particularly significant in enhancing the quality of life for CI recipients, allowing them to navigate and engage with their surroundings more effectively.
I-A Related work
@@
Many reviews have been written in the context of CI. For example, [11] discussed the advantages offered by machine learning (ML) to cochlear implantation, such as analyzing data to personalize treatment strategies. It enhances accuracy in speech processing optimization, surgical anatomy location prediction, and electrode placement discrimination. Besides, it delves into its applications, including optimizing cochlear implant fitting, predicting patient threshold levels, and automating image-guided CI surgery. The review discusses some novel opportunities for research, emphasizing the need for high-quality data inputs and addressing concerns about algorithm transparency in clinical decision-making for improved patient care. Similarly, the review by Manero et al. [12] details some benefits of employing artificial intelligent (AI) in enhancing CI technology, involving adaptive sound processing, acoustic scene classification, and auditory scene analysis. The authors discuss AI-driven advancements aiming to optimize sound signals, adapt to diverse environments, and improve speech perception for individuals with hearing loss, ultimately enhancing their overall quality of life.
Additionally, the review [13] explores three main topics: direct-speech neuroprosthesis, which involves decoding speech from the sensorimotor cortex using AI, including the synthesis of produced speech from brain activity; a top-down exploration of pediatric cochlear implantation using ML, delving into its applications in pediatric cochlear implantation; and the potential of AI to solve the hearing-in-noise problem, examining its capabilities in addressing challenges related to hearing in noisy environments. Moreover, the review [14] critically examines the current landscape of tele-audiology practices, highlighting both their constraints and potential opportunities. Specifically, it explores intervention and rehabilitation efforts for CIs, focusing on remote programming and the concept of self-fitting CIs. Recently, a review by Henry et al. in 2023 [15] conducts a comprehensive review of noise reduction algorithms employed in CIs. Maintaining a general classification based on the number of microphones used—single or multiple channels—the analysis extends to incorporate recent studies showcasing a growing interest in ML techniques. The review culminates with an exploration of potential research avenues that hold promise for future advancements in the field. Table I offers a comparative analysis of the proposed review in contrast to other discussed AI-based CI reviews and surveys.
@@
TABLE I: Comparison of this review against other existing AI-based CI reviews and surveys. Tick marks (✓) indicate that a particular field has been considered, while cross marks () indicate that a field has been left unaddressed. The symbol (★) indicates that the most critical concerns of a field have not been addressed.
Ref. Year Description Backg. TCI Metrics datasets MLT DLT Apps RGCs FDs
[11] 2019 The intersection of ML and CIs
[12] 2022 Explores materials and devices designed for medical care
[13] 2022 AI in otolaryngology and communication sciences
[14] 2022 Remote audiology services
[15] 2023 Signal processing in CI for noise reduction
Our 2024 Advanced AI algorithms for CIs
@@
I-B Statistics on investigated papers
Recently, there has been a surge in publications related to AI-based CI. The review methodology entails defining the search strategy and study selection criteria. Criteria for inclusion, such as keyword relevance and impact, shape the quality assessment protocol. A comprehensive search was conducted on databases such as Scopus and Web of Science. Keywords were extracted for theme clustering, resulting in a formulated query to gather advanced AI-based CI studies. The research query retrieves references from papers containing the keywords ”Cochlear implant” or ”Hearing loss” and ”Artificial intelligence” in their abstracts, titles, or authors’ keywords. It subsequently refines these papers, focusing on those that also include ”Machine learning,” ”Deep learning,” or ”Reinforcement learning.” Figure 1 illustrates the most frequently used keywords by the authors in the titles, abstracts, and keywords of the selected papers.
@@ I-C Motivation and contribution
  • The implementation of CI, along with a comprehensive elucidation of the taxonomy encompassing ML and DL-based CIs, is thoroughly expounded upon. Additionally, recommended frameworks for AI-based CI are thoroughly discussed and succinctly summarized in tables for enhanced clarity.
@@
  • Delving into the existing gaps in AI-driven CI, offering insights, and proposing novel ideas to address these gaps. Additionally, exploring potential avenues for future research to deepen comprehension and provide valuable guidance for subsequent investigations.
The subsequent sections are organized as follows: Section II delves into the background in speech processing, outlining datasets and metrics. Section III discusses the methodology employed for CI based on AI. Section IV presents the medical applications and impact of applying AI on CI. Section V offers a comprehensive discussion on research gaps, future directions, and perspectives. Finally, Section VI concludes the paper with implications and future research directions.
II Background in speech processing for CIs
II-A Speech processing for CIs
CIs are electronic devices that can be implanted in one or both ears to restore some level of hearing for individuals with partial or severe deafness. CIs comprise an external part with a microphone and speech processor and an internal part with a receiver-stimulator and electrode array, as shown in Figure 3. They convert sounds into electrical signals, stimulating the auditory nerve to enable sound perception in individuals with profound hearing loss [16].
@@ II-A Speech processing for CIs
The incoming sound is divided into multiple frequency channels using bandpass filters and then processed by envelope detectors. Non-linear compressors adjust the dynamic range of the envelope for each patient. The compressed envelope amplitudes are then utilized to modulate a fixed-rate biphasic carrier signal. A current source converts voltage into pulse trains of current, which are delivered to electrodes placed along the cochlea in a non-overlapping manner. This stimulation method is called continuous interleaved sampling (CIS). Another coding strategy, known as advanced combination encoder (ACE), uses a greater number of channels and dynamically selects the ”n-of-m” bands with the largest envelope amplitudes (prior to compression). Only the corresponding ”n” electrodes are stimulated. A popular device widely used for CIs, such as the cochlear nucleus, typically has 22 channels.
The sound processor, which usually contains a microphone, battery, and other components, can be worn either behind-the-ear (BTE) or off-the-ear (OTE). A headpiece holds a transmitter coil, positioned externally above the ear, while internally, a receiver coil, stimulator, and electrode array are implanted. The SP includes a digital signal processor (DSP) with memory units (maps) that store patient-specific information. An audiologist configures these maps during the fitting process, adjusting thresholds for each electrode, including T-Levels (the softest current levels audible to the CI user) and C/M-Levels (current levels perceived as comfortably loud), as well as the stimulation rate or programming strategy. Data (pulse amplitude, pulse duration, pulse gap, etc.) and power are sent through the skull via a radio frequency signal from the transmitter coil to the receiver coil. The stimulator decodes the received bitstream and converts it into electric currents to be delivered to the cochlear electrodes. High-frequency signals stimulate electrodes near the base of the cochlea, while low-frequency signals stimulate electrodes near the apex.
@@
The CI stimulates the auditory nerve afferents, which connect to the central auditory pathways. However, compared to individuals with normal hearing (NH), CI users face more difficulties in speech perception, particularly in noisy environments. Hearing loss can be caused by various factors, including natural aging, genetic predisposition, exposure to loud sounds, and medical treatments. Damage to the hair cells in the inner ear often leads to a reduced dynamic range of hearing, as well as decreased frequency selectivity and discriminative ability in speech processing. To evaluate the effectiveness of CIs in speech perception amid noise, listening tests involving both normal hearing individuals and CI users are commonly conducted. These tests typically employ a combination of speech utterances from a recognized speech corpus and background noises like speech-weighted noise and babble. Alternatively, vocoder simulations can be utilized alongside speech intelligibility metrics. While sentence-based tests are frequently employed, other stimuli such as vowels, consonants, and phonemes are also used. As a result, noise reduction techniques are increasingly employed to enhance the performance of CIs in challenging environments [15].
II-B Datasets
@@
Researchers have utilized numerous datasets to validate their proposed schemes, comprising both widely recognized publicly available sets and locally generated ones. These datasets fall into two categories: speech or images. Table II provides a summary of these datasets, detailing their characteristics, citing studies that have utilized them, and indicating their availability through links or references.
TABLE II: List of publicly available datasets used for CIs applications
@@
PhonDat 1 21587 201 It is from the Bavarian Archive For Speech Signals, contains 21.4 hours of speech data, and it is orthographically transcribed and phonemically annotated. [21],[22] Link [^1] https://www.phonetik.uni-muenchen.de/Bas/BasPD1eng.html TIMIT 6300 630 The dataset consists of phonemically and lexically transcribed speech from American English speakers belonging to diverse demographics and dialects. It provides comprehensive information with time-aligned orthographic, phonetic, and word transcriptions. Additionally, each utterance is accompanied by its corresponding 16-bit, 16kHz speech waveform file, ensuring a complete and detailed dataset for analysis and experimentation in ASR and acoustic-phonetic studies. [23],[24],[25] Link [^2] https://www.kaggle.com/mfekadu/darpa-timit-acousticphonetic-continuous-speech GSC 18 Hours 30 The dataset was gathered using crowd-sourcing. It consists of 65,000 recordings, each lasting one second, and contains 30 brief words. Among these words, 20 commonly used ones were spoken five times by the majority of participants, while 10 other words (considered unfamiliar) were spoken only once. [26] Link [^3] https://research.googleblog.com/2017/08/launching-speech-commands-dataset.html WSJ0-2mix - - This dataset is a collection of mixed speech recordings that are used for speech recognition. It utilizes utterances from the Wall Street Journal (WSJ0) corpus to create these mixtures. [27] Link  [^4] https://www.merl.com/research/highlights/deep-clustering LibriVox 547 Hours - The LibriVox dataset consists of sentence-aligned German audio, text, and English translations from audio books. [24] Link  [^5] https://librivox.org/ HSM 20 30 The development of the German HSM Sentence Test aimed to provide an ample amount of test sentences for the repeated assessment of speech comprehension among CI users. [24],[28] [29] CEC1 6000 24 The dataset features simulated living rooms with static sources, including a single target speaker, interferer (competing talker or noise), and a large target speech database of English sentences produced by 40 British English speakers. [28] Link [^6] https://claritychallenge.org/clarity_CEC1_doc/docs/cec1_data DEMAND 560 6 The dataset consists of 15 recordings capturing acoustic noise in various environments. These recordings were made using a 16-channel array, with microphone distances ranging from 5 cm to 21.8 cm. [24] Link [^7] https://www.kaggle.com/datasets/chrisfilo/demand THCHS-30 35 Hours 50 The dataset is a free Chinese speech corpus accompanied by resources such as lexicon and language models. [30] Link [^8] https://www.openslr.org/18/ BCP 55938 20 The bern cocktail party (BCP) dataset contains Cocktail Party scenarios with individuals wearing CI audio processors and a head and torso simulator. Recorded in an acoustic chamber, it includes multi-channel audio, image recordings, and digitized microphone positions for each participant. [31] [32] ikala 252-30s 206 The dataset comprises audio recordings consisting of vocal and backing track music with a sampling rate of 44100 Hz. Each music track is a stereo recording, where one channel contains the singing voice and the other channel contains the background music. All tracks were performed by professional musicians and featured a group of six singers, evenly split between three females and three males. [33] Link [^9] https://zenodo.org/records/3532214 MUSDB 150 4 The dataset is a collection of music tracks specifically designed for music source separation research. It consists of professionally mixed songs across various genres, with individual tracks isolated for vocals, drums, bass, and other accompaniment. [33] Link [^10] https://sigsep.github.io/datasets/musdb.html CQ500 491 - The dataset includes anonymized dicoms files, along with the interpretations provided by radiologists. The interpretations were conducted by three radiologists who have 8, 12, and 20 years of experience in interpreting cranial CT scans, respectively. [34] Link [^11] http://headctstudy.qure.ai/dataset ECochG 21 - The dataset captures inner ear potentials in response to acoustic stimulation, specifically focusing on cochlear implant recipients. [35],[36],[37] Link [^12] https://zenodo.org/records/7092661 LibriSpeech 1000 Hours 2484 speakers English speech recordings with sampling rate 16kHz, the recordings are sourced from audiobooks read aloud for the LibriVox dataset and have been meticulously segmented and synchronized. [38] Link [^13] https://www.openslr.org/12 WHAM! 81.68 Hours 28000 files It contains pairs two-speaker mixtures with unique noise backgrounds. WHAMR! extends this by adding artificial reverberation. Noise audio, collected from various urban locations, is provided with scripts to build the datasets. [38] Link [^14] https://wham.whisper.ai/
@@
II-C Metrics
Multiple evaluation metrics are utilized during the training and validation of any DL models, including AI-based CI. These metrics, integral to the confusion matrix, are widely known and applicable across various data types such as speech or images. They include Accuracy (Acc), Sensitivity (Sen), Recall (Rec), Specificity (Spe), Precision (Pre), F1 score (F1), and receiver operating characteristic curve (ROC). Moreover, different metrics play roles in prediction tasks. For instance, intersection over union (IoU) assesses overlap, and mean absolute error (MAE) quantifies absolute differences. For a comprehensive understanding of the metrics discussed, including their equations, refer to the details provided in [5, 39]. Other metrics that are widely used for CI are summarized in Table LABEL:tab:metrics.
TABLE III: An overview of the metrics employed for evaluating CI methods.
Metric Formula Description
DCS $\frac{{2\cdot|A\cap B|}}{{|A|+|B|}}$ The metric dice coefficient similarity (DCS) is used to evaluate the performance of the vestibule segmentation network. The Dice coefficient is a widely used similarity metric in image segmentation tasks. It measures the overlap between the predicted segmentation mask and the ground truth mask.
ASD $\frac{\sum_{a\in S(A)}{\min_{b\in S(B)}(a-b)}+\sum_{b\in S(B)}{\min_{a\in S(A)}(b-a)}}{|S(A)|+|S(B)|}$Ac ASD is a commonly used evaluation measure in medical image segmentation tasks. It quantifies the average distance between the surfaces of two segmented objects, typically the predicted segmentation and the ground truth.
AVD $\max(\text{d}(A,B),\text{d}(B,A))$Ac AVD quantifies the average difference in volume between a predicted segmentation and a reference or ground truth segmentation..
SNR $\mathrm{10\log_{10}\Big{(}\frac{Signal}{Noise}\Big{)}_{dB}}$Ac SNR is a measure of the quality of the speech signal. It is commonly used to evaluate the quality of the stego-speech (the speech signal after the hidden information has been embedded). A lower SNR indicates that the steganography technique has introduced more distortion to the speech signal.
ASE $\frac{\sum E_{ij}}{N}$ Is the average surface error (ASE). The distances between corresponding points on the measured surface and the reference surface are computed using $E_{ij}=|I_{1}(i,j)-I_{2}(k,l)|$, known as point to point error (P2PE). $N$ is the total number of correspondence points. Normalize average error by dividing by the imaging system’s dynamic range. $(i,j)$, and $(k,l)$ represent pixel values in the first and second images, respectively.
ILD $20\log_{10}\left(\frac{L}{R}\right)$Ac ILD a psychoacoustic metric, measuring sound level differences between left (L) and right (R) ears, reflects cues essential for sound localization, enhancing spatial awareness in auditory perception for directional sound source identification.
MEE $\frac{1}{N}\sum_{i=1}^{N}|L_{i}-\hat{L}_{i}|$Ac MEE measures the average absolute difference between the true $L_{i}$ and estimated $\hat{L}_{i}$ endpoint locations across multiple utterances $N$.
MMSE Estimator $\hat{x}_{MMSE}=E[X|Y]$Ac MMSE is a statistical estimation technique used in speech enhancement to minimize the mean square error between the estimated $Y$ and true $X$ clean speech signals
STOI $\frac{\sum_{t=1}^{T}g(t)\cdot\text{STOI}_{\text{frame}}(t)}{\sum_{t=1}^{T}g(t)}$Ac STOI is a metric used to assess the intelligibility of time-frequency weighted noisy speech. It is based on the idea that human speech perception relies on the availability of important acoustic features in short time frames [40].